# Jone Lopez de Gamiz Zearra Stevenson Center 1323

### v2

@@ 1. Introduction
One of the consequences of the characterization by Gordon and Wise is that coherent Artin groups whose defining graph is not complete can be decomposed as a free amalgamated product along a free abelian subgroup. We consider more general Artin groups admitting such a decomposition and show
@@
In the last section, we turn to a more geometric perspective and consider the notion of acylindrical hyperbolicity (see Section 5). There are many results on the literature about acylindrical hyperbolicity of various Artin groups, and a summary can be found in [8] where Charney, Martin and Morris-Wright conjecture that if $A_{\Gamma}$ is an Artin group that does not decompose as a direct product of standard parabolic subgroups, then the quotient $A_{\Gamma}/Z(A_{\Gamma})$ is acylindrically hyperbolic. Moreover, they prove this conjecture for even Artin groups of FC-type. Here, we consider acylindrical hyperbolicity for subgroups of even Artin groups of FC-type.
Even Artin groups are Artin groups for which all the labels in the defining graph are even numbers. They obviously include RAAGs and share some of their nice properties. For example, as for RAAGs, subgroups associated to subgraphs (called standard parabolic subgroups) are retracts. Even Artin groups of FC-type can be defined as even Artin groups with the condition that in any triangle in the defining graph there are always at least two edges labeled by $2$. They are even closer to RAAGs; for instance, Antolín and Foniqi [1] have shown that, as in the class of RAAGs, an arbitrary intersection of parabolic subgroups in an even Artin group of FC-type is again parabolic, and the same property has been conjectured for arbitrary Artin groups (see [8]). Antolín and Foniqi’s result is essential for our next main theorem, which generalizes an analogous theorem for RAAGs by Minasyan and Osin [21].
@@
Finally, we use this theorem to obtain the following corollary. Note that an acylindrically hyperbolic group cannot contain two normal non-trivial subgroups that intersect trivially (see [21, Theorem 3.7]). Thus, this result reflects that this obstruction yields a characterization of acylindrical hyperbolicity for subgroups of even Artin groups of FC-type.
@@ 2.1. Artin groups
We will abuse notation and write $T\subseteq\Gamma$ if $T$ is a subgraph of $\Gamma$. We say that a subgraph $T\subseteq\Gamma$ is induced if whenever two vertices of $T$ are linked in $\Gamma$, then they are also linked in $T$ with the same label. Then, by [5], the subgroup of $W_{\Gamma}$ generated by the vertices of $T$ is isomorphic to the Coxeter group $W_{T}$, and analogously, by [24], the subgroup of $A_{\Gamma}$ generated by the vertices of $T$ is isomorphic to the Artin group $A_{T}$ (and we will use $A_{T}$ to denote this subgroup). Also, if $S,T\subseteq\Gamma$, then as subgroups of $A_{\Gamma}$ we have that $A_{S}\cap A_{T}=A_{S\cap T}$ (see [24]).
Remark 2.3.
@@
Note that if $A_{\Gamma}$ is spherical, then the graph $\Gamma$ must be complete.
@@
The conjecture is known to be true in several particular cases, including the cases of spherical Artin groups, RAAGs and, more generally, Artin groups of FC-type. In addition, an Artin group $A_{\Gamma}$ is called 2-dimensional if for any triangle subgraph $\Delta\subseteq\Gamma$, the Artin group $A_{\Delta}$ is not spherical. The spherical triangle Artin groups are precisely those whose associated graph has labels $(2,2,k),(2,3,3),(2,3,4)$ or $(2,3,5)$. The $K(\pi,1)$-Conjecture is also true for $2$-dimensional Artin groups, and this implies that the standard presentation complex is a $K(\pi,1)$, so in particular, it is aspherical.
Another interesting property of even Artin groups is that standard parabolic subgroups are retracts. More explicitly, if $T\subseteq\Gamma$ is a subgraph, then there is a well-defined epimorphism
@@ 2.2. Coherent right-angled Artin groups
[11, Theorem 1] If $A_{\Gamma}$ is a RAAG, then it is coherent if and only if $\Gamma$ is chordal, that is, if $\Gamma$ does not contain a subgraph isomorphic to a cycle of length greater than $3$.
@@
[14, Theorem 2] If $A_{\Gamma}$ is a RAAG, then it is coherent if and only if its derived subgroup $[A_{\Gamma},A_{\Gamma}]$ is free.
@@ 3. Coherent Artin groups
If there is a vertex $t$ in $\Gamma_{u}$ adjacent to $w$, then there is a cycle through $v,w,t$ and $u$. From condition (i), we deduce that there is some vertex $a$ in $\Gamma_{2}\setminus\Delta$ linked to both $v$ and $w$, and condition (ii) forces these two labels to be equal to $2$.
@@
For arbitrary RAAGs, it is well known that the abelianization of a standard parabolic subgroup coincides with its intersection with the abelianization of the ambient group. We first extend this fact to the class of even Artin groups.
@@
Let $\Delta$ be the complete labeled simplicial graph with the same number of vertices as $\Gamma$ and such that all the labels are $2$. That is, $A_{\Delta}$ is the abelianization of $A_{\Gamma}$ and if $\varphi$ is the abelianization map $\varphi\colon A_{\Gamma}\to A_{\Delta}$, then $[A_{\Gamma},A_{\Gamma}]=\ker\varphi$. But the fact that $A_{\Gamma}$ is even implies that the restriction of $\varphi$ to $A_{\Omega}$ is also the abelianization map of $A_{\Omega}$, so
@@
We furthermore show that if $A_{\Gamma}$ is a coherent Artin group with the decomposition as in Proposition 3.2, then the conclusion of Proposition 3.5 still remains true for the standard parabolic subgroups $A_{\Gamma_{1}},A_{\Gamma_{2}}$ and $A_{\Delta}$.
@@
The inclusion $[A_{\Omega},A_{\Omega}]\subseteq A_{\Omega}\cap[A_{\Gamma},A_{\Gamma}]$ is obvious. For the converse, let $Z$ be the complete labeled simplicial graph with the same number of vertices as connected components in $\Gamma^{odd}$ and such that all the labels are $2$. For each vertex $v$ of $\Gamma$, let $\varphi(v)$ be the vertex in $Z$ corresponding to the connected component of $\Gamma^{odd}$ that contains $v$. Then there is a well-defined group homomorphism
$\varphi\colon A_{\Gamma}$ $\to A_{Z}$
$v$ $\mapsto\varphi(v).$
Indeed, note that every defining relation in $A_{\Gamma}$ corresponds to an edge in $\Gamma$ between two vertices, say $v$ and $w$. If $\varphi(v)=\varphi(w)$ then the relation is obviously preserved, and this is always the case if the label is odd. If $\varphi(v)\neq\varphi(w)$, then the label must be even and, by the definition of $Z$, $\varphi(v)\varphi(w)=\varphi(w)\varphi(v)$. Hence, the relation is also preserved in this case.
Moreover, $A_{Z}$ is precisely the abelianization of the group $A_{\Gamma}$, so $\varphi$ is the abelianization map and $[A_{\Gamma},A_{\Gamma}]=\ker\varphi$. It remains to show that $\varphi_{|A_{\Omega}}$ is the abelianization map of the standard parabolic subgroup $A_{\Omega}$. For that aim, it suffices to show that the connected components of $\Omega^{odd}$ are precisely the intersections of the connected components of $\Gamma^{odd}$ with $\Omega$.
@@
[20] Let $G=G_{1}\ast_{H}G_{2}$ be a free amalgamated product and let $N\trianglelefteq G$ be a normal subgroup such that $N\cap H=1$ and both intersections $N\cap G_{1}$, $N\cap G_{2}$ are free. Then the subgroup $N$ is free.
At this point we have all the necessary tools to prove one of the implications of Theorem 3.3.
@@
Let $A_{\Gamma}$ be a coherent Artin group. Then the derived subgroup $[A_{\Gamma},A_{\Gamma}]$ is free.
@@
If $A_{\Gamma}$ is a coherent Artin group, from Proposition 3.2 we know that either $A_{\Gamma}$ is the direct product of a dihedral Artin group and a free abelian group or it has a decomposition as in the second bullet point. In the first case, the result is trivial, so let us assume that $A_{\Gamma}$ decomposes as
@@
Lemma 3.6 implies that $[A_{\Gamma},A_{\Gamma}]\cap A_{\Delta}=[A_{\Delta},A_{\Delta}]=1$. Moreover, by induction we may suppose that $[A_{\Gamma},A_{\Gamma}]\cap A_{\Gamma_{1}}$ and $[A_{\Gamma},A_{\Gamma}]\cap A_{\Gamma_{2}}$ are free groups. In conclusion, from Lemma 3.8 we deduce that the derived subgroup $[A_{\Gamma},A_{\Gamma}]$ is also free. ∎
@@
Therefore, $H_{2}(N)\otimes_{G/N}F\neq 0$, which implies that $H_{2}(N)\neq 0$. Thus, the group $N$ is not free and, moreover, $H_{2}(N)$ cannot be a finitely generated abelian group because, in that case, it would be a torsion $\mathbb{Z}[G/N]$-module. ∎
@@
If $\Gamma$ is as in (iii), we distinguish two cases depending on whether $A_{\Gamma}$ is spherical or not. The spherical cases are precisely when the labels of the triangle are $(2,3,3)$, $(2,3,4)$ or $(2,3,5)$ (see, for instance, [19]). In the $(2,3,4)$ case, from [15, Corollary 1.5] we see that the derived subgroup is finitely generated but not finitely presented, so it is not free. In the other two remaining cases, both groups have infinite cyclic abelianizations. If the derived subgroup was free, we would have that
@@
This means that $A_{\Gamma}$ is $2$-dimensional (for non spherical triangle Artin groups see [18, Lemma 2.2]), so $A_{\Gamma}$ satisfies the $K(\pi,1)$-Conjecture. Hence, the presentation complex of the standard presentation is aspherical. The fact that $\Gamma$ is a cycle implies that if $n$ is the number of vertices, then this presentation has precisely $n$ generators and $n$ relations, so from Lemma 3.10 we deduce that the derived subgroup is not free. ∎
@@ 4. Finitely generated normal subgroups of coherent Artin groups
where $\Gamma_{1},\Gamma_{2},\Delta$ are proper subgraphs, $\Delta=\Gamma_{1}\cap\Gamma_{2}$ and $A_{\Delta}$ is a free abelian group. Denote by $T$ the Bass-Serre tree associated to this decomposition. Then the kernel $K$ of the action of $G$ on $T$ is a standard parabolic subgroup of $G$ which is a free abelian direct factor. That is, $K=A_{S}$ where $S$ is a subgraph of $\Gamma$, $K$ is a free abelian group and
@@
Let us now prove the claim. Note that we may actually assume that $G$ is generated by $w$ and $\Delta$, that is, we may assume that $\Gamma=\Delta\cup\{w\}$ and that $S$ equals $\operatorname{lk}^{2}_{\Gamma}(w)\cap\Delta$. In this setting, we may split the groups as follows:
@@
At this point, we are going to use the result conjectured by Tits and proved by Crisp-Paris in [9, Theorem 1] stating that, in an Artin group, the squares of the standard generators generate a RAAG with relators being the commutators between those vertices that already commuted in the ambient group. More precisely, if we consider the subgroup $C_{1}$ of $C=A_{\Delta}$ generated by $t^{2}$ for $t\in\Delta$ and the subgroup $A_{1}$ of $A$ generated by $w^{2}$ and $C_{1}$, then by [9, Theorem 1], $A_{1}$ is the free product of the cyclic group generated by $w^{2}$ and the free abelian group $C_{1}$. Thus $C_{1}$ is malnormal in $A_{1}$, so in particular $C_{1}\cap C_{1}^{w^{2}}=1$. As the subgroups $C_{1}$ and $C_{1}^{w^{2}}$ have finite index in $C$ and $C^{w^{2}}$, respectively, $1=C_{1}\cap C_{1}^{w^{2}}$ has finite index in $A_{\Delta}\cap A_{\Delta}^{w^{2}}$, so $A_{\Delta}\cap A_{\Delta}^{w^{2}}=1$.
@@
The group $N$ is a subgroup of $A_{\Gamma}$, so it also acts on $T$ and by [22, Theorem 2.1] we have that either $NA_{\Delta}$ has finite index in $A_{\Gamma}$ or $N$ lies in $A_{\Delta}$. In the former case, since $A_{\Delta}$ is a free abelian group, we get that $A_{\Gamma}/N$ is virtually abelian. In the later case, if $N\trianglelefteq A_{\Delta}$, as $N$ is normal, it actually lies in $K$, the kernel of the action of $A_{\Gamma}$ on $T$, so the result follows from Lemma 4.2. ∎

### v3

# Jone Lopez de Gamiz Zearra Facultad de Economía y Empresa, Bilbao-Sarriko Universidad del País Vasco EHU/UPV

@@ 1. Introduction
Before stating some other results proved in the article, we would like to point out that in the paper, by a subgraph we will mean induced subgraph (see Definition 2.3).
One of the consequences of the characterization by Gordon and Wise is that coherent Artin groups whose defining graph is not complete can be decomposed as a free amalgamated product along a free abelian subgroup. We consider more general Artin groups admitting such a decomposition in terms of proper subgraphs and we get the following.
@@
In the last section, we turn to a more geometric perspective and consider the notion of acylindrical hyperbolicity (see Section 5). There are many results on the literature about acylindrical hyperbolicity of various Artin groups, and a summary can be found in [10]. In [22], based on questions posted in [9], Haettel conjectures that if $A_{\Gamma}$ is an Artin group that does not decompose as a direct product of standard parabolic subgroups, then the quotient $A_{\Gamma}/Z(A_{\Gamma})$ is acylindrically hyperbolic. Moreover, they prove this conjecture for even Artin groups of FC-type. Here, we consider acylindrical hyperbolicity for subgroups of even Artin groups of FC-type.
Even Artin groups are Artin groups for which all the labels in the defining graph are even numbers. They obviously include RAAGs and share some of their nice properties. For example, as for RAAGs, subgroups associated to subgraphs (called standard parabolic subgroups) are retracts (see Subsetcion 2.1).
Even Artin groups of FC-type can be defined as even Artin groups with the condition that in any triangle in the defining graph there are always at least two edges labeled by $2$. They are even closer to RAAGs; for instance, Antolín and Foniqi [1] have shown that, as in the class of RAAGs, an arbitrary intersection of parabolic subgroups in an even Artin group of FC-type is again parabolic, and the same property has been conjectured for arbitrary Artin groups (see [10]). Antolín and Foniqi’s result is essential for our next main theorem, which generalizes an analogous theorem for RAAGs by Minasyan and Osin [25].
@@
Finally, we use this theorem to obtain the following corollary. Note that an acylindrically hyperbolic group cannot contain two infinite normal subgroups that intersect trivially (see [25, Theorem 3.7]). Furthermore, even Artin groups of FC-type satisfy the $K(\pi,1)$-Conjecture (see Section 2.1), so they are torsion-free. Thus, Theorem 5.3 reflects that this obstruction yields a characterization of acylindrical hyperbolicity for subgroups of even Artin groups of FC-type.
@@
The first named author has been supported by the Spanish Government grant PID2020-117281GB-I00, partly by the European Regional Development Fund (ERDF), and the Basque Government, grant IT1483-22. The second named author has been partially supported by the Spanish Government PID2021-126254NB-I00 and Departamento de Ciencia, Universidad y Sociedad del Conocimiento del Gobierno de Aragón (grant code: E22-23R: “Álgebra y Geometría”). We thank the anonymous referees for their comments and suggestions that helped to improve the paper.
@@ 2.1. Artin groups
We will abuse notation and write $T\subseteq\Gamma$ if $T$ is a subgraph of $\Gamma$.
Definition 2.3.
We say that a subgraph $T\subseteq\Gamma$ is induced if whenever two vertices of $T$ are linked in $\Gamma$, then they are also linked in $T$ with the same label.
@@
If $T\subseteq\Gamma$ is a subgraph (which in our context means induced), by [5], the subgroup of $W_{\Gamma}$ generated by the vertices of $T$ is isomorphic to the Coxeter group $W_{T}$, and analogously, by [30], the subgroup of $A_{\Gamma}$ generated by the vertices of $T$ is isomorphic to the Artin group $A_{T}$ (and we will use $A_{T}$ to denote this subgroup). Also, if $S,T\subseteq\Gamma$, then as subgroups of $A_{\Gamma}$ we have that $A_{S}\cap A_{T}=A_{S\cap T}$ (see [30]).
@@
Note that if $A_{\Gamma}$ is spherical, then the graph $\Gamma$ must be complete. A particular type of spherical Artin groups are the so called dihedral Artin groups which correspond to the case when the graph $\Gamma$ consists of two vertices joined by an edge.
@@
The conjecture is known to be true in several particular cases, including the cases of spherical Artin groups, RAAGs and, more generally, Artin groups of FC-type. In addition, an Artin group $A_{\Gamma}$ is called 2-dimensional if for any triangle subgraph $\Delta\subseteq\Gamma$, the Artin group $A_{\Delta}$ is not spherical. The spherical triangle Artin groups are precisely those whose associated graph has labels $(2,2,k),(2,3,3),(2,3,4)$ or $(2,3,5)$. The $K(\pi,1)$-Conjecture is also true for $2$-dimensional Artin groups.
An equivalent formulation of the $K(\pi,1)$-Conjecture is that certain complex constructed by Salvetti, known as the Salvetti complex, is a model for $K(\pi,1)$, [8]. In the case of a 2-dimensional Artin group this complex is precisely the standard presentation complex, so from the fact that 2-dimensional Artin groups satisfy the $K(\pi,1)$-Conjecture, it follows that the presentation complex for these groups is aspherical.
Another interesting and well-known property of even Artin groups is that standard parabolic subgroups are retracts (see, for example, [1]). More explicitly, if $T\subseteq\Gamma$ is a subgraph, then there is a well-defined epimorphism
@@ 2.2. Coherent right-angled Artin groups
[13, Theorem 1] If $A_{\Gamma}$ is a RAAG, then it is coherent if and only if $\Gamma$ is chordal, that is, if $\Gamma$ does not contain a subgraph isomorphic to a cycle of length greater than $3$.
@@
[16, Theorem 2] If $A_{\Gamma}$ is a RAAG, then it is coherent if and only if its derived subgroup $[A_{\Gamma},A_{\Gamma}]$ is free.
@@ 3. Coherent Artin groups
If there is a vertex $t$ in $\Gamma_{u}$ adjacent to $w$, then there is a cycle through $v,w,t$ and $u$. We claim that there is some vertex $a$ in $\Gamma_{2}\setminus\Delta$ linked to both $v$ and $w$. Condition (i) implies that there is a chord in the cycle through $v,w,t$ and $u$. That chord yields a shorter cycle through $v,w$ and $\Gamma_{u}$, so repeating this argument periodically, we eventually get the claim. Finally, condition (ii) forces the labels between $a$ and $v$ and between $a$ and $w$ to be equal to $2$.
@@
For arbitrary RAAGs, it is well-known that the abelianization of a standard parabolic subgroup coincides with its intersection with the abelianization of the ambient group (see [16, Section 4]). We first extend this fact to the class of even Artin groups.
@@
Let $\Delta$ be the complete labeled simplicial graph with the same number of vertices as $\Gamma$ and such that all the labels are $2$. As we are assuming that $A_{\Gamma}$ is even, there is a well-defined epimorphism $\varphi\colon A_{\Gamma}\to A_{\Delta}$ which is precisely the abelianization map. Therefore $[A_{\Gamma},A_{\Gamma}]=\ker\varphi$. But using again the fact that $A_{\Gamma}$ is even we deduce that the restriction of $\varphi$ to $A_{\Omega}$ is also the abelianization map of $A_{\Omega}$, so
@@
We furthermore show that if $A_{\Gamma}$ is a coherent Artin group with the decomposition as in Proposition 3.2, then the conclusion of Proposition 3.5 still remains true for the standard parabolic subgroups $A_{\Gamma_{1}},A_{\Gamma_{2}}$ and $A_{\Delta}$. But before that, it will be useful to state the following (probably well-known) result that describes the abelianization of an arbitrary Artin group.
Lemma 3.6.
Let $A_{\Gamma}$ be an Artin group and let $Z$ be the complete labeled simplicial graph with the same number of vertices as connected components in $\Gamma^{odd}$ and such that all the labels are $2$. Then
$A_{\Gamma}/[A_{\Gamma},A_{\Gamma}]\cong A_{Z}.$
Proof.
For each vertex $v$ of $\Gamma$, let $\varphi(v)$ be the vertex in $Z$ corresponding to the connected component of $\Gamma^{odd}$ that contains $v$. Then there is a well-defined group homomorphism
$\varphi\colon A_{\Gamma}$ $\to A_{Z}$
$v$ $\mapsto\varphi(v).$
Indeed, note that every defining relation in $A_{\Gamma}$ corresponds to an edge in $\Gamma$ between two vertices, say $v$ and $w$. If $\varphi(v)=\varphi(w)$, then the relation is obviously preserved, and this is always the case if the label is odd. If $\varphi(v)\neq\varphi(w)$, then the label must be even and, by the definition of $Z$, $\varphi(v)\varphi(w)=\varphi(w)\varphi(v)$. Hence, the relation is also preserved in this case. Moreover, $\varphi$ is obviously surjective. Given any map $\psi\colon A_{\Gamma}\to A$ where $A$ is an abelian group, if $v$ and $w$ are vertices in $\Gamma$ linked by an edge with odd label, the fact that $A$ is abelian implies that $\psi(v)=\psi(w)$. Hence, the same holds true for all the vertices in the same connected component of $\Gamma^{odd}$, and this means that $\psi$ factors through $\varphi$. ∎
@@
The inclusion $[A_{\Omega},A_{\Omega}]\subseteq A_{\Omega}\cap[A_{\Gamma},A_{\Gamma}]$ is obvious. Consider the map $\varphi$ defined in Lemma 3.6. The proof of that result implies that $\varphi$ is the abelianization map and that $[A_{\Gamma},A_{\Gamma}]=\ker\varphi$. It remains to show that $\varphi_{|A_{\Omega}}$ is the abelianization map of the standard parabolic subgroup $A_{\Omega}$. For that aim, it suffices to show that the connected components of $\Omega^{odd}$ are precisely the intersections of the connected components of $\Gamma^{odd}$ with $\Omega$.
@@
[24] Let $G=G_{1}\ast_{H}G_{2}$ be a free amalgamated product and let $1\neq N\trianglelefteq G$ be a normal subgroup such that $N\cap H=1$ and both intersections $N\cap G_{1}$, $N\cap G_{2}$ are free or trivial. Then the subgroup $N$ is free.
At this point we have all the necessary tools to prove one of the implications of Theorem 3.3. We begin with the special case of dihedral Artin groups.
Example 3.10.
Dihedral Artin groups are well-known to be free-by-cyclic: in fact this is also a way to see that they are coherent [19]. Being free-by-cyclic implies that the derived subgroup is free. This is also a consequence of the explicit computations performed on [26, Theorems 3.6, 3.9 and 3.18].
@@
Let $A_{\Gamma}$ be a coherent Artin group. Then the derived subgroup $[A_{\Gamma},A_{\Gamma}]$ is free or trivial.
@@
If $A_{\Gamma}$ is abelian, $[A_{\Gamma},A_{\Gamma}]$ is trivial. If $A_{\Gamma}$ is a non abelian coherent Artin group, from Proposition 3.2 we know that either $A_{\Gamma}$ is the direct product of a dihedral Artin group and a free abelian group or it has a decomposition as in the second bullet point. In the first case, the result is trivial by Example 3.10.
In the second case, let us assume that $A_{\Gamma}$ decomposes as
@@
Lemma 3.7 implies that $[A_{\Gamma},A_{\Gamma}]\cap A_{\Delta}=[A_{\Delta},A_{\Delta}]=1$. Moreover, by induction on the number of vertices we may suppose that $[A_{\Gamma_{1}},A_{\Gamma_{1}}]=[A_{\Gamma},A_{\Gamma}]\cap A_{\Gamma_{1}}$ and $[A_{\Gamma_{2}},A_{\Gamma_{2}}]=[A_{\Gamma},A_{\Gamma}]\cap A_{\Gamma_{2}}$ are either free or trivial groups. In conclusion, from Lemma 3.9 we deduce that the derived subgroup $[A_{\Gamma},A_{\Gamma}]$ is also free. ∎
@@
Therefore, $H_{2}(N)\otimes_{G/N}F\neq 0$, which implies that $H_{2}(N)\neq 0$. Thus, the group $N$ is not free and, moreover, $H_{2}(N)$ cannot be a finitely generated abelian group because, in that case, it would be a torsion $\mathbb{Z}[G/N]$-module. This implies that $N$ is not finitely presented. Otherwise, one could construct a free resolution with finitely generated modules up to degree 2, which would force $H_{2}(N)$ to be finitely generated. ∎
@@
If $\Gamma$ is as in (iii), we distinguish two cases depending on whether $A_{\Gamma}$ is spherical or not. The spherical cases are precisely when the labels of the triangle are $(2,3,3)$, $(2,3,4)$ or $(2,3,5)$ (see, for instance, [23]). In the $(2,3,4)$ case, from [17, Corollary 1.5] we see that the derived subgroup is finitely generated but not finitely presented, so it is not free. In the other two remaining cases, both groups have infinite cyclic abelianizations. If the derived subgroup was free, we would have that
@@
This means that $A_{\Gamma}$ is $2$-dimensional (for non spherical triangle Artin groups see [21, Lemma 2.2]), so $A_{\Gamma}$ satisfies the $K(\pi,1)$-Conjecture. Hence, the presentation complex of the standard presentation is aspherical. The fact that $\Gamma$ is a cycle implies that if $n$ is the number of vertices, then this presentation has precisely $n$ generators and $n$ relations, so from Lemma 3.12 we deduce that the derived subgroup is not free (observe that Lemma 3.6 implies that the abelianization of $A_{\Gamma}$ is torsion-free). ∎
@@ 4. Finitely generated normal subgroups of coherent Artin groups
where $\Gamma_{1},\Gamma_{2},\Delta$ are proper subgraphs, $\Delta=\Gamma_{1}\cap\Gamma_{2}$ and $A_{\Delta}$ is a free abelian group. Denote by $T$ the Bass-Serre tree associated to this decomposition. Then the kernel $K$ of the action of $A_{\Gamma}$ on $T$ is a standard parabolic subgroup of $A_{\Gamma}$ which is a free abelian direct factor. That is, $K=A_{S}$ where $S$ is a subgraph of $\Gamma$, $K$ is a free abelian group and
@@
Let us now prove the claim. Note that we may actually assume that $A_{\Gamma}$ is generated by $w$ and $\Delta$, that is, we may assume that $\Gamma=\Delta\cup\{w\}$ and that $S$ equals $\operatorname{lk}^{2}_{\Gamma}(w)\cap\Delta$. In this setting, we may split the groups as follows:
@@
At this point, we are going to use the result conjectured by Tits and proved by Crisp-Paris in [11, Theorem 1] stating that, in an Artin group, the squares of the standard generators generate a RAAG with relators being the commutators between those vertices that already commuted in the ambient group. More precisely, if we consider the subgroup $C_{1}$ of $C=A_{\Delta}$ generated by $t^{2}$ for $t\in\Delta$ and the subgroup $A_{1}$ of $A$ generated by $w^{2}$ and $C_{1}$, then by [11, Theorem 1], $A_{1}$ is the free product of the cyclic group generated by $w^{2}$ and the free abelian group $C_{1}$. Thus $C_{1}$ is malnormal in $A_{1}$, so in particular $C_{1}\cap C_{1}^{w^{2}}=1$. As the subgroups $C_{1}$ and $C_{1}^{w^{2}}$ have finite index in $C$ and $C^{w^{2}}$, respectively, $1=C_{1}\cap C_{1}^{w^{2}}$ has finite index in $A_{\Delta}\cap A_{\Delta}^{w^{2}}$, so $A_{\Delta}\cap A_{\Delta}^{w^{2}}=1$.
@@
The group $N$ is a subgroup of $A_{\Gamma}$, so it also acts on $T$ and by [27, Theorem 2.1] we have that either $NA_{\Delta}$ has finite index in $A_{\Gamma}$ or $N$ lies in $A_{\Delta}$. In the former case, since $A_{\Delta}$ is a free abelian group, we get that $A_{\Gamma}/N$ is virtually abelian. In the later case, if $N\trianglelefteq A_{\Delta}$, as $N$ is normal, it actually lies in $K$, the kernel of the action of $A_{\Gamma}$ on $T$, so the result follows from Lemma 4.2. ∎